OpenStudy (anonymous):

1/2t+17/6=55/18 solve using the addition principle add or the multiplication principle..the solution to t=

OpenStudy (anonymous):

\[\Large \frac12 \cdot t+\frac{17}{6}=\frac{55}{18}\] \[\Large \frac t2=\frac{55}{18}-\frac{17}{6}\] .. go on continue...

OpenStudy (anonymous):

simplity the asnewr

OpenStudy (anonymous):

thats not mulitification \[1/2t + 17/6-55/18 solve\]

OpenStudy (anonymous):

\[\Large \frac t2= \frac{55\cdot 6-17\cdot 18}{18\cdot 6}\] \[\Large \frac t2= \frac{330-306}{108}\] \[\Large \frac t2=\frac{24}{108}\] \[\Large \frac t2=\frac29\] \[\LARGE 9t=4 \longrightarrow \boxed{t=\cfrac94}\]

OpenStudy (anonymous):

t=4/9

OpenStudy (anonymous):

not 9/4

OpenStudy (anonymous):

9=6+7/3t

OpenStudy (anonymous):

new ?